OpenStudy (anonymous):

A stone is thrown straight up from the edge of a roof, 675 feet above the ground, at a speed of 18 feet per second. A. Remembering that the acceleration due to gravity is -32 feet per second squared, how high is the stone 6 seconds later? B. At what time does the stone hit the ground? C. What is the velocity of the stone when it hits the ground?

OpenStudy (nincompoop):

this is easy

OpenStudy (anonymous):

i need help setting it up

OpenStudy (anonymous):

i understand anti derivatives and finding constants… setting up the problem is the hard part

OpenStudy (nincompoop):

we can set this up using integration if you want

OpenStudy (nincompoop):

let us begin with the -32 ft/sec

OpenStudy (anonymous):

that would be v(t)?

OpenStudy (nincompoop):

it's up to you, I would set this up using s(t) you can use v(t) or a(t) it doesn't matter

OpenStudy (anonymous):

ok, just tell me how you would do it ;)

OpenStudy (nincompoop):

list the following -32 as your rate of fall 18 as your initial velocity 675 as your initial position

OpenStudy (nincompoop):

integrate -32 in terms of time for me

OpenStudy (nincompoop):

don't you think I know what I am doing, annas?

OpenStudy (anonymous):

physics nazi!! haha its ok i understand

OpenStudy (nincompoop):

fine fine s^2 geez

OpenStudy (anonymous):

so… s(t) = 16x^2

OpenStudy (anonymous):

srry 16t^2

OpenStudy (nincompoop):

damn... you're skipping steps

OpenStudy (anonymous):

v(t) = 32t

OpenStudy (anonymous):

a(t) = 32

OpenStudy (nincompoop):

\[\int\limits -32 dt=32t+C\]

OpenStudy (anonymous):

constants! right..

OpenStudy (anonymous):

so v(18) = 32(18) + C1 = -32(18)?

OpenStudy (anonymous):

so c = 18/16

OpenStudy (anonymous):

-18/16

OpenStudy (nincompoop):

huh? don't you put zero where the t's are?

OpenStudy (anonymous):

wait, wouldn't it be (-32t + c) though?

OpenStudy (nincompoop):

\[s'(t) = \int\limits s''(t) = \int\limits\limits\limits -32dt=-16t+C \rightarrow s'(0)=-16t+C=18\]

OpenStudy (nincompoop):

\[ s'(0)=-16t+C=18 \rightarrow -16(0) + C = 18 \rightarrow C= 18\]

OpenStudy (anonymous):

so a(t) = -32 v(t) = -16t + c1 s(t) = -8t^2 +c1(t) +c2 thats all i understand

OpenStudy (nincompoop):

okay, we can do that you wanna change from a to v to s

OpenStudy (anonymous):

yes plz

OpenStudy (nincompoop):

so we just solved your c1 as 18

OpenStudy (nincompoop):

v(t) = -16t + 18

OpenStudy (nincompoop):

then let us integrate v(t) to find s(t)

OpenStudy (anonymous):

so s(t) = -8t^2 + 18t +C2

OpenStudy (nincompoop):

\[s(t) = \int\limits v(t) = \int\limits (-16t+18) dt = \int\limits -16t dt + \int\limits 18dt \]

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

lead coefficient should still be negative, yes?

OpenStudy (nincompoop):

\[s(t) = \frac{ -16t^2 }{ 2 }+ 18t + C_2 \rightarrow s(0) = -8t^2+18t+C_2=675\]

OpenStudy (anonymous):

so c2 = 675. got it

OpenStudy (nincompoop):

\[C_2 = 675 \rightarrow s(t) = -8t^2 +18t+675\]

OpenStudy (nincompoop):

so that is your position function

OpenStudy (anonymous):

s(t) = -8t^2 + 18t + 675 A). s(6) = -8(6)^2 + 18(6) + 675

OpenStudy (nincompoop):

yeah pretty much

OpenStudy (nincompoop):

math, skull

OpenStudy (nincompoop):

applications of integration

OpenStudy (anonymous):

why isn't this answer correct :(

OpenStudy (nincompoop):

who said it isn't?

OpenStudy (nincompoop):

it say how HIGH, so it is asking for the position at t= 6

OpenStudy (nincompoop):

then at what time does the stone hit the ground now your position function should be equals to zero then solve for t

OpenStudy (anonymous):

ok i got it :) thank you!!

OpenStudy (nincompoop):

s(6) = -8(6)^2 + 18(6) + 675 = 495 ft

OpenStudy (nincompoop):

s(t) = -8t^2 + 18t + 675 = 0 use QUADRATIC EQ or any tricks you know to solve for t

OpenStudy (nincompoop):

wanna do it together?

OpenStudy (anonymous):

i got B now

OpenStudy (anonymous):

how do you get C?

OpenStudy (nincompoop):

t = -8.1292 and t = 10.379 pick the positive

OpenStudy (nincompoop):

well, how do you find the velocity when s(t) = 0?

OpenStudy (nincompoop):

you use the velocity v(t) that means you take the derivative of s(t) but since you already know that you just use it

OpenStudy (anonymous):

would i plug v(10.379)?

OpenStudy (nincompoop):

ya ya

OpenStudy (nincompoop):

-8(10.379)^2 + 18(10.379) = -674.967128

OpenStudy (anonymous):

boom! 100%

OpenStudy (anonymous):

thank you!

OpenStudy (nincompoop):

therefore it is -674.967128 ft/s^2

OpenStudy (anonymous):

XOXO

OpenStudy (nincompoop):

that stone is going to turn into diamond

OpenStudy (nincompoop):

piece of cake, yeh?

OpenStudy (nincompoop):

are we still doing the same problem?

OpenStudy (nincompoop):

you can close it now